INFRARED SPECTROSCOPY OF TWO ISOMERS OF THE OCS-CS_{2} COMPLEX

\underline{J. N. OLIAEE}, M. DEHGHANY, MAHIN AFSHARI, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, 2500 University Dr., N.W., Calgary, Alberta T2N 1N4, Canada; A.R.W. McKELLAR, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario, K1A 0R6, Canada.

A second isomer of OCS-CS_{2} complex, with a nonplanar cross-shaped structure, has been studied for the first time by analysing an infrared band in the region of the OCS stretching vibration (2062 cm^{−1}). This isomer has C_{s} symmetry and the observed band consists of purely c-type rotational transitions. The ground state rotational parameters are found to be A=0.07306 cm^{−1}, B=0.03325 cm^{−1} and C=0.02879 cm^{−1}, in good agreement with a previous semi-empirical calculation.^{a} In addition, a hybrid band with a- and b-type rotational transitions has been assigned to be due to the known planar form of OCS-CS_{2}, as previously studied by microwave spectroscopy.^{a} The spectra were recorded using a rapid-scan tunable diode laser spectrometer to probe a pulsed supersonic jet expansion. Calculations indicate that the planar isomer of OCS-CS_{2} is the lowest in energy, in contrast to CO_{2}-CS_{2} where the crossed form is believed to be the lowest.^{b}

aJ. J. Newby, M. M. Serafin, R. A. Peebles and S. A. Peebles, Phys. Chem. Chem. Phys. 7, 487 (2005).

bC. C. Dutton, D. A. Dows, R. Eikey, S. Evans and R. A. Beaudet, J. Phys. Chem. A 102, 6904 (1998).